The current in a certain conductor is 40mA

No 1. Find the total charge in coulombs that passes through to the conductor in 1.5s

No.2. Find the total number of electrons that pass through the conductor in that same time


1
Expert's answer
2021-07-06T08:22:24-0400

1.- We have to consider that I = 40 mA = 0.040 A and then we find the charge with the time t=1.5 s by using "Q=It" :


"Q=It=(0.040\,A)(1.5\,s)=(0.040\frac{C}{\cancel{s}})(1.5\,\cancel{s})"


"Q=0.06\text{ Coulomb}"


2.- If we remember that each electron has a charge equal to "q_e=1.602\times10^{-19}C" we can define the relationships between the total charge Q and qe to find the number of electrons N:


"Q=Nq_e \implies N=\frac{Q}{q_e}=\frac{0.06\,\cancel{C}}{1.602\times10^{-19}\cancel{C}/electron}=\frac{0.06}{1.602}\times10^{19}\text{ electrons}"


"N=3.745\times10^{17}\text{ electrons}"


In conclusion, (No 1.) the total charge that goes through the conductor in such conditions is 0.06 C (No. 2) and approximately 3.745 X 1017 electrons traveled through the conductor as well in that same time period.


Reference:

  • Young, H. D., & Freedman, R. A. (2015). University Physics with Modern Physics and Mastering Physics (p. 1632). Academic Imports Sweden AB.